OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 15m √9m and width √3m, which of the following is the area of a rectangle?

OpenStudy (anonymous):

A. 45m^2√3 B. 45m√3m C.135m^2√3 D.135m√3m

OpenStudy (akashdeepdeb):

Everything is given to you. All you need to do is multiply them. You need to know one thing for this though... √a * √b = √ab [understood?] So √8 * √6 = √8*6 = √48 So here area of the rectangle will be Length times width = l*w = 15√9 * √3 = 15√27 = 15*√3*3*3 = 15*3√3 = 45√3 Understood? :)

OpenStudy (akashdeepdeb):

Is that m or meters? If it is m then area = 15m√9m * √3m = 15m√27m^2 = 15m^2*3√3 = 45m^2√3 Understood? :)

OpenStudy (anonymous):

mhmmm :)

OpenStudy (akashdeepdeb):

Congratulations: This is my 700th Math question :)

OpenStudy (akashdeepdeb):

:D You should be proud. ://// :P

OpenStudy (anonymous):

Hahah yay!

OpenStudy (akashdeepdeb):

Please like this video and my above reply :p http://www.youtube.com/watch?v=yVPSvYDvN04

OpenStudy (anonymous):

Simplify: \[\frac{ x ^{2}-3x-18 }{ x ^{2}-36 }\]